OpenStudy (anonymous):

1. y = –2x + 5 (1 point)

OpenStudy (shaik0124):

what to do

OpenStudy (anonymous):

what you want??

OpenStudy (anonymous):

1. y = –2x + 5 (1 point)

OpenStudy (shaik0124):

does u require value of x?

OpenStudy (anonymous):

x =2

OpenStudy (shaik0124):

y=1

OpenStudy (shaik0124):

y=1 is answer if x=2

OpenStudy (anonymous):

okthank you can you answer 2 more for I really need this

OpenStudy (anonymous):

2. y = 7x (1 point) 3. y = 3x – 9 (1 point)

OpenStudy (shaik0124):

y=7*2=14 y=3*2-9=6-9=-3 y=-3 is answer for 2 nd one

OpenStudy (shaik0124):

@muhammadjack got it

OpenStudy (shaik0124):

@muhammadjack

OpenStudy (anonymous):

ok whats the answer for 3

OpenStudy (shaik0124):

for 2nd is y=14 for 3rd y=-3

OpenStudy (anonymous):

thank you very much you just saved my life from a lot of pain